# Hardware Acceleration for Neural Networks: A Comprehensive Survey

### v1

@@ Abstract.
Neural networks have become a dominant computational workload across cloud and edge platforms, but their rapid growth in model size and deployment diversity has exposed hardware bottlenecks that are increasingly dominated by memory movement, communication, and irregular operators rather than peak arithmetic throughput. This survey reviews the current technology landscape for hardware acceleration of deep learning, spanning GPUs and tensor-core architectures, domain-specific accelerators (e.g., TPUs/NPUs), FPGA-based designs, ASIC inference engines, and emerging LLM-serving accelerators such as LPUs (language processing units), alongside in-/near-memory computing and neuromorphic/analog approaches. We organize the space using a unified taxonomy across (i) workloads (CNNs, RNNs, GNNs, Transformers/LLMs), (ii) execution settings (training vs. inference; datacenter vs. edge), and (iii) optimization levers (reduced precision, sparsity and pruning, operator fusion, compilation and scheduling, and memory-system/interconnect design). We synthesize key architectural ideas such as systolic arrays, vector and SIMD engines, specialized attention and softmax kernels, quantization-aware datapaths, and high-bandwidth memory, and we discuss how software stacks and compilers bridge model semantics to hardware. Finally, we highlight open challenges—including efficient long-context LLM inference (KV-cache management), robust support for dynamic and sparse workloads, energy- and security-aware deployment, and fair benchmarking—pointing to promising directions for the next generation of neural acceleration.
@@ 1. Introduction
Figure 1. Overview of hardware acceleration
Figure 1 provides a high-level overview of the hardware-acceleration landscape covered in this survey.
@@
Deep neural networks now underpin a wide range of applications—from perception and control in autonomous and cyber-physical systems to search, recommendation, and natural language assistants (Krizhevsky et al., 2012; Vaswani et al., 2017; Xu et al., 2025c). At the same time, modern models have expanded rapidly in scale (parameters, activation sizes, and context lengths), diversity (CNNs, GNNs, Transformers, diffusion models), and deployment environments (datacenter training clusters, latency-critical edge inference, and battery-powered on-device intelligence) (Abadi et al., 2016; Gunter et al., 2024; Xu et al., 2025d, b). These trends have elevated hardware acceleration from a performance optimization to an enabling technology: without specialized compute, memory systems, and software co-design, many state-of-the-art models would be impractical due to cost, latency, or energy constraints (Hennessy and Patterson, 2019; Jouppi et al., 2017).
@@
Why acceleration is harder than “more FLOPs” is that, although matrix multiplication remains central, end-to-end efficiency is increasingly governed by data movement rather than arithmetic (Williams et al., 2009). Training and inference pipelines stress memory capacity and bandwidth (e.g., activation storage and optimizer state during training; KV-cache and batching dynamics during LLM inference) (Micikevicius et al., 2018; Kwon et al., 2023), interconnect bandwidth and collectives for distributed execution, and irregular operators arising from sparsity, dynamic control flow, and mixture-of-experts routing (Gale et al., 2019; Shazeer et al., 2017). As a result, accelerator design must address the full system stack: compute datapaths, memory hierarchies, on-chip networks, off-chip DRAM/HBM, host-device interfaces, and multi-device interconnects, together with compilers and runtime systems that map high-level models onto hardware efficiently (Abadi et al., 2016).
@@
The technology landscape spans general-purpose GPUs enhanced with tensor cores and high-bandwidth memory; domain-specific processors such as TPUs and mobile/edge NPUs (Jouppi et al., 2017, 2021); FPGA-based accelerators offering reconfigurability and tight integration with custom I/O pipelines (Nurvitadhi et al., 2017; Umuroglu et al., 2017; Venieris and Bouganis, 2017); ASIC inference engines optimized for throughput-per-watt (Chen et al., 2014); and LLM-serving accelerators such as LPUs (language processing units) that target predictable low-latency token generation (Groq, 2024). Beyond digital CMOS datapaths, in-/near-memory computing and analog approaches attempt to reduce the fundamental cost of memory movement (Shafiee et al., 2016), while neuromorphic and event-driven designs explore alternative execution models for sparse, low-power workloads (Davies et al., 2018). In practice, the most capable systems emerge from hardware–software co-design: model architectures and training recipes that are amenable to quantization and sparsity (Jacob et al., 2018; Han et al., 2016b); compilers that perform fusion, layout transformations, and scheduling; and kernels and dataflows that exploit tensorization, tiling, and locality (Kung, 1982; Williams et al., 2009).
@@
In this survey, we focus on current technology and architectural principles that generalize across devices and workloads. We structure the discussion along three axes: (i) workloads (CNNs, RNNs, GNNs, and Transformers; and key operators such as attention, convolution, normalization, and sampling) (Vaswani et al., 2017), (ii) execution settings (training vs. inference; offline vs. online; cloud vs. edge), and (iii) optimization levers (reduced precision, structured and unstructured sparsity, compression and pruning, operator fusion, compilation and scheduling, and memory/interconnect design) (Micikevicius et al., 2018; Jacob et al., 2018; Gale et al., 2019; Han et al., 2016b; Williams et al., 2009). Rather than treating hardware in isolation, we highlight how these levers interact—for example, how low precision shifts bottlenecks toward memory, how sparsity challenges utilization and load balance, and how long-context inference amplifies memory-capacity constraints (Kwon et al., 2023).
@@ 2. Accelerator Challenges
Figure 2. Limitations of hardware acceleration
Figure 2 summarizes key limitations that shape accelerator design choices across performance, energy, and deployability.
Designing and deploying neural-network accelerators requires navigating a multi-objective optimization problem where improvements along one axis often degrade another. In practice, “speed” is not a single number: training emphasizes time-to-train and scalability, while inference emphasizes latency, tail latency under load, and throughput at a target quality. Across both settings, energy and resource budgets constrain what can be deployed, especially when models must run continuously or at the edge (Hennessy and Patterson, 2019).
@@ 2.1. Power and energy consumption
Figure 3. Power consumption
Figure 3 highlights the power-consumption perspective that motivates energy-aware mappings and reduced data movement.
Precision and sparsity are primary levers for energy reduction. Mixed-precision training improves throughput and reduces memory traffic while maintaining convergence, typically using reduced-precision compute with higher-precision accumulation and stabilization techniques (Micikevicius et al., 2018). For inference, quantization to integer arithmetic can significantly reduce energy and improve effective bandwidth, but it introduces calibration and accuracy trade-offs that depend on the model and deployment target (Jacob et al., 2018). Compression and sparsity reduce memory footprint and traffic, but unstructured sparsity can add indexing overhead and reduce utilization if the hardware/software stack is not sparsity-aware (Han et al., 2016b; Gale et al., 2019).
@@ 2.1.2. TPU/NPU-specific energy challenges
Tensor processors improve efficiency by specializing dense dataflows and maximizing on-chip reuse, but their energy advantage diminishes when workloads deviate from the assumed shapes or require frequent off-chip traffic (Jouppi et al., 2017, 2021; Williams et al., 2009). Attention-heavy workloads and long-context serving stress memory systems via KV-cache and intermediate activations, pushing energy back toward DRAM/HBM movement (Vaswani et al., 2017; Kwon et al., 2023; Horowitz, 2014). In edge NPUs, power is constrained by SoC-level thermal limits and shared memory bandwidth; as a result, sustained performance-per-watt depends on compiler scheduling, operator coverage, and quantized datapaths (Jacob et al., 2018; Mazumder et al., 2021). When sparsity or conditional execution is present (e.g., MoE), energy efficiency can suffer due to reduced reuse and load imbalance unless the stack explicitly supports these patterns (Shazeer et al., 2017; Gale et al., 2019).
@@ 2.1.3. ASIC-specific energy challenges
ASICs can minimize energy by hardwiring data reuse, choosing an energy-efficient dataflow, and provisioning large on-chip SRAM to reduce DRAM access (Chen et al., 2016, 2014; Horowitz, 2014). However, energy efficiency can degrade when models include unsupported operators (e.g., attention variants or custom normalization) that force host fallback or additional data marshaling (Vaswani et al., 2017; Chen et al., 2016). Compression and sparsity reduce weight/activation traffic, but the benefits depend on sparsity structure; irregular sparsity introduces indexing overhead and load imbalance that can negate energy savings (Han et al., 2016a, b; Gale et al., 2019). Designs closer to sensors (e.g., vision-centric accelerators) reduce I/O energy, yet they are specialized and may not generalize to broader model families (Liu et al., 2015).
@@ 2.1.4. FPGA-specific energy challenges
FPGAs can be energy-efficient when a streaming pipeline eliminates intermediate DRAM writes and uses custom precision to reduce switching activity and memory traffic (Umuroglu et al., 2017; Ma et al., 2018; Horowitz, 2014). In practice, performance-per-watt is sensitive to external memory bandwidth, on-chip BRAM/DSP utilization, and routing overhead; poor mapping can waste energy through underutilized pipelines and excess off-chip transfers (Nurvitadhi et al., 2017; Venieris and Bouganis, 2017). Transformer/attention-style workloads further increase bandwidth pressure and can reduce the energy advantage if buffering and fusion are insufficient (Zhang et al., 2024a; Vaswani et al., 2017). In cloud FPGA deployments, host-device interfaces and multi-tenant overheads can dominate system energy unless amortized by sustained utilization and carefully engineered serving pipelines (Fowers et al., 2018; Nurvitadhi et al., 2017).
@@ 2.1.6. In-/near-memory and analog energy challenges
Analog in-/near-memory approaches reduce data movement for matmul-heavy layers by performing multiply-accumulate in crossbars, but ADC/DAC overheads, calibration, and device non-idealities can offset energy benefits at the system level (Shafiee et al., 2016; Chi et al., 2016; Ankit et al., 2019; Horowitz, 2014). Energy efficiency is sensitive to mapping: if only a fraction of the model maps to crossbars, the overall system can still be dominated by digital compute and data conversion costs (Ankit et al., 2019). Attention-heavy models introduce softmax and KV-cache access patterns that are not naturally accelerated by crossbar matmul and can erode end-to-end savings (Vaswani et al., 2017; Kwon et al., 2023). Consequently, practical designs often require heterogeneous integration and careful partitioning between analog and digital components (Hennessy and Patterson, 2019).
@@ 2.1.7. Neuromorphic energy challenges
Neuromorphic systems can achieve extremely low energy when activity is sparse and event-driven computation matches the workload, but benefits are sensitive to encoding/decoding overheads and to whether sparsity is realized in practice (Davies et al., 2018; Merolla et al., 2014). When spike rates increase, communication and routing energy rises, and the event-driven advantage diminishes (Davies et al., 2018). Mapping conventional dense networks to spiking representations can add conversion overhead and reduce effective sparsity, making comparisons with dense accelerators sensitive to modeling assumptions (Gale et al., 2019; Coleman et al., 2019). As a result, neuromorphic energy wins are most robust in always-on sensing and temporal workloads where sparse events are intrinsic to the data (Merolla et al., 2014).
@@ 2.2. Throughput, latency, and speed
Figure 4. LLM inference bottleneck
For LLM inference, the critical path often alternates between compute-heavy projection/MLP GEMMs and bandwidth-heavy attention/KV-cache reads, as illustrated in Figure 4. IO-aware attention kernels reduce memory traffic by tiling attention and avoiding materialization of large intermediate tensors, improving both throughput and latency under memory pressure (Dao et al., 2022). Runtime techniques such as paging KV-cache help maintain throughput at high concurrency, but they introduce their own trade-offs in fragmentation, scheduling overhead, and tail latency (Kwon et al., 2023).
At the software level, kernel fusion and code generation reduce launch overhead and improve locality, which matters especially for small GEMMs and latency-sensitive inference. Libraries and compiler stacks (e.g., vendor primitives, graph compilers, and kernel DSLs) play a central role in realizing these optimizations in practice (Chetlur et al., 2014; Chen et al., 2018; Tillet et al., 2019). Consequently, performance claims should specify not only the hardware but also the kernel/library and compilation path used.

### v3

@@ Abstract.
Neural networks have become a dominant computational workload across cloud and edge platforms, but their rapid growth in model size and deployment diversity has exposed hardware bottlenecks that are increasingly dominated by memory movement, communication, and irregular operators rather than peak arithmetic throughput. This survey reviews the current technology landscape for hardware acceleration of deep learning, spanning Graphics Processing Units (GPUs) and tensor-core architectures, domain-specific accelerators (e.g., Tensor Processing Units (TPUs)/Neural Processing Units (NPUs)), Field-Programmable Gate Array (FPGA)-based designs, Application-Specific Integrated Circuit (ASIC) inference engines, and emerging Large Language Model (LLM)-serving accelerators such as Language Processing Units (LPUs), alongside in-/near-memory computing and neuromorphic/analog approaches. We organize the survey using a unified taxonomy across (i) workloads (Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), Transformers/Large Language Models (LLMs)), (ii) execution settings (training vs. inference; datacenter vs. edge), and (iii) optimization levers (reduced precision, sparsity and pruning, operator fusion, compilation and scheduling, and memory-system/interconnect design). We synthesize key architectural ideas such as systolic arrays, vector and Single Instruction, Multiple Data (SIMD) engines, specialized attention and softmax kernels, quantization-aware datapaths, and high-bandwidth memory, and we discuss how software stacks and compilers bridge model semantics to hardware. Finally, we highlight open challenges—including efficient long-context LLM inference (Key-Value (KV)-cache management), robust support for dynamic and sparse workloads, energy- and security-aware deployment, and fair benchmarking—pointing to promising directions for the next generation of neural acceleration.
@@ 1. Introduction
Figure 1. High-level overview of the hardware acceleration landscape, illustrating the spectrum from general-purpose GPUs to domain-specific TPUs, FPGAs, and ASICs, along with the interaction between compute datapaths, memory hierarchies, and the software stack.
@@
Deep neural networks now underpin a wide range of applications—from perception and control in autonomous and cyber-physical systems (Xu et al., 2025d, a) to search, recommendation, and natural language assistants (Krizhevsky et al., 2012; Vaswani et al., 2017). At the same time, modern models have expanded rapidly in scale (parameters, activation sizes, and context lengths), diversity (Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), Transformers, diffusion models), and deployment environments (datacenter training clusters, latency-critical edge inference, and battery-powered on-device intelligence) (Abadi et al., 2016; Gunter et al., 2024). These trends have elevated hardware acceleration from a performance optimization to an enabling technology: without specialized compute, memory systems, and software co-design, many state-of-the-art models would be impractical due to cost, latency, or energy constraints (Hennessy and Patterson, 2019; Jouppi et al., 2017).
@@
Why acceleration is harder than “more Floating Point Operations (FLOPs)” is that, although matrix multiplication remains central, end-to-end efficiency is increasingly governed by data movement rather than arithmetic (Williams et al., 2009). Training and inference pipelines stress memory capacity and bandwidth (e.g., activation storage and optimizer state during training; Key-Value (KV)-cache and batching dynamics during Large Language Model (LLM) inference) (Micikevicius et al., 2018; Kwon et al., 2023), interconnect bandwidth and collectives for distributed execution, and irregular operators arising from sparsity, dynamic control flow, and Mixture-of-Experts (MoE) routing (Gale et al., 2019; Shazeer et al., 2017). As a result, accelerator design must address the full system stack: compute datapaths, memory hierarchies, on-chip networks, off-chip Dynamic Random Access Memory (DRAM)/High-Bandwidth Memory (HBM), host-device interfaces, and multi-device interconnects, together with compilers and runtime systems that map high-level models onto hardware efficiently (Abadi et al., 2016).
@@
The technology landscape shown in Figure 1 spans general-purpose Graphics Processing Units (GPUs) enhanced with tensor cores and high-bandwidth memory; domain-specific processors such as Tensor Processing Units (TPUs) and mobile/edge Neural Processing Units (NPUs) (Jouppi et al., 2017, 2021); Field-Programmable Gate Array (FPGA)-based accelerators offering reconfigurability and tight integration with custom I/O pipelines (Nurvitadhi et al., 2017; Umuroglu et al., 2017; Venieris and Bouganis, 2016; Boutros et al., 2024; Yan et al., 2024); Application-Specific Integrated Circuit (ASIC) inference engines optimized for throughput-per-watt (Chen et al., 2014); and LLM-serving accelerators such as Language Processing Units (LPUs) that target predictable low-latency token generation (Groq, 2024). Beyond digital Complementary Metal-Oxide-Semiconductor (CMOS) datapaths, in-/near-memory computing and analog approaches attempt to reduce the fundamental cost of memory movement (Shafiee et al., 2016), while neuromorphic and event-driven designs explore alternative execution models for sparse, low-power workloads (Davies et al., 2018). In practice, the most capable systems emerge from hardware–software co-design: model architectures and training recipes that are amenable to quantization and sparsity (Jacob et al., 2018; Han et al., 2016b); compilers that perform fusion, layout transformations, and scheduling; and kernels and dataflows that exploit tensorization, tiling, and locality (Kung, 1982; Williams et al., 2009).
@@
In this survey, we focus on current technology and architectural principles that generalize across devices and workloads. We structure the discussion along three axes: (i) workloads (CNNs, Recurrent Neural Networks (RNNs), GNNs, and Transformers; and key operators such as attention, convolution, normalization, and sampling) (Vaswani et al., 2017), (ii) execution settings (training vs. inference; offline vs. online; cloud vs. edge), and (iii) optimization levers (reduced precision, structured and unstructured sparsity, compression and pruning, operator fusion, compilation and scheduling, and memory/interconnect design) (Micikevicius et al., 2018; Jacob et al., 2018; Gale et al., 2019; Han et al., 2016b; Williams et al., 2009; Wang et al., 2024; Li et al., 2024b; Fang et al., 2024; Muralidharan et al., 2024). Rather than treating hardware in isolation, we highlight how these levers interact—for example, how low precision shifts bottlenecks toward memory, how sparsity challenges utilization and load balance, and how long-context inference amplifies memory-capacity constraints (Kwon et al., 2023).
@@ 2. Accelerator Challenges
Figure 2. Key limitations and trade-offs in hardware acceleration design, highlighting the multi-objective optimization problem involving performance (throughput/latency), energy efficiency, and deployment constraints such as area and cost.
Figure 2 summarizes key limitations that shape accelerator design choices across performance, energy, and deployability. Designing and deploying neural-network accelerators requires navigating a multi-objective optimization problem where improvements along one axis often degrade another. In practice, “speed” is not a single number: training emphasizes time-to-train and scalability, while inference emphasizes latency, tail latency under load, and throughput at a target quality. Across both settings, energy and resource budgets constrain what can be deployed, especially when models must run continuously or at the edge (Hennessy and Patterson, 2019).
@@ 2.1. Power and energy consumption
Figure 3. Power consumption analysis highlighting (a) the dominance of data movement energy over arithmetic operations, (b) the impact of reduced precision on energy efficiency, and (c) the trade-offs involved in exploiting sparsity.
Figure 3 highlights the power-consumption perspective that motivates energy-aware mappings and reduced data movement. Precision and sparsity are primary levers for energy reduction. Mixed-precision training improves throughput and reduces memory traffic while maintaining convergence, typically using reduced-precision compute with higher-precision accumulation and stabilization techniques (Micikevicius et al., 2018). For inference, quantization to integer arithmetic can significantly reduce energy and improve effective bandwidth, but it introduces calibration and accuracy trade-offs that depend on the model and deployment target (Jacob et al., 2018). Compression and sparsity reduce memory footprint and traffic, but unstructured sparsity can add indexing overhead and reduce utilization if the hardware/software stack is not sparsity-aware (Han et al., 2016b; Gale et al., 2019).
@@ 2.1.2. TPU/NPU-specific energy challenges
Tensor processors improve efficiency by specializing dense dataflows and maximizing on-chip reuse, but their energy advantage diminishes when workloads deviate from the assumed shapes or require frequent off-chip traffic (Jouppi et al., 2017, 2021; Williams et al., 2009). Attention-heavy workloads and long-context serving stress memory systems via KV-cache and intermediate activations, pushing energy back toward DRAM/HBM movement (Vaswani et al., 2017; Kwon et al., 2023; Horowitz, 2014). In edge NPUs, power is constrained by System-on-Chip (SoC)-level thermal limits and shared memory bandwidth; as a result, sustained performance-per-watt depends on compiler scheduling, operator coverage, and quantized datapaths (Jacob et al., 2018; Mazumder et al., 2021). When sparsity or conditional execution is present (e.g., MoE), energy efficiency can suffer due to reduced reuse and load imbalance unless the stack explicitly supports these patterns (Shazeer et al., 2017; Gale et al., 2019).
@@ 2.1.3. ASIC-specific energy challenges
ASICs can minimize energy by hardwiring data reuse, choosing an energy-efficient dataflow, and provisioning large on-chip Static Random Access Memory (SRAM) to reduce DRAM access (Chen et al., 2016, 2014; Horowitz, 2014). However, energy efficiency can degrade when models include unsupported operators (e.g., attention variants or custom normalization) that force host fallback or additional data marshaling (Vaswani et al., 2017; Chen et al., 2016). Compression and sparsity reduce weight/activation traffic, but the benefits depend on sparsity structure; irregular sparsity introduces indexing overhead and load imbalance that can negate energy savings (Han et al., 2016a, b; Gale et al., 2019). Designs closer to sensors (e.g., vision-centric accelerators) reduce I/O energy, yet they are specialized and may not generalize to broader model families (Liu et al., 2015).
@@ 2.1.4. FPGA-specific energy challenges
FPGAs can be energy-efficient when a streaming pipeline eliminates intermediate DRAM writes and uses custom precision to reduce switching activity and memory traffic (Umuroglu et al., 2017; Ma et al., 2018; Horowitz, 2014). In practice, performance-per-watt is sensitive to external memory bandwidth, on-chip Block RAM (BRAM)/Digital Signal Processor (DSP) utilization, and routing overhead; poor mapping can waste energy through underutilized pipelines and excess off-chip transfers (Nurvitadhi et al., 2017; Venieris and Bouganis, 2016). Transformer/attention-style workloads further increase bandwidth pressure and can reduce the energy advantage if buffering and fusion are insufficient (Zhang et al., 2024a; Vaswani et al., 2017). In cloud FPGA deployments, host-device interfaces and multi-tenant overheads can dominate system energy unless amortized by sustained utilization and carefully engineered serving pipelines (Fowers et al., 2018; Nurvitadhi et al., 2017).
@@ 2.1.6. In-/near-memory and analog energy challenges
Analog in-/near-memory approaches reduce data movement for matmul-heavy layers by performing multiply-accumulate in crossbars, but Analog-to-Digital Converter (ADC)/Digital-to-Analog Converter (DAC) overheads, calibration, and device non-idealities can offset energy benefits at the system level (Shafiee et al., 2016; Chi et al., 2016; Ankit et al., 2019; Horowitz, 2014). Energy efficiency is sensitive to mapping: if only a fraction of the model maps to crossbars, the overall system can still be dominated by digital compute and data conversion costs (Ankit et al., 2019). Attention-heavy models introduce softmax and KV-cache access patterns that are not naturally accelerated by crossbar matmul and can erode end-to-end savings (Vaswani et al., 2017; Kwon et al., 2023). Consequently, practical designs often require heterogeneous integration and careful partitioning between analog and digital components (Hennessy and Patterson, 2019).
@@ 2.1.7. Neuromorphic energy challenges
Neuromorphic systems can achieve extremely low energy when activity is sparse and event-driven computation matches the workload, but benefits are sensitive to encoding/decoding overheads and to whether sparsity is realized in practice (Davies et al., 2018; Merolla et al., 2014). When spike rates increase, communication and routing energy rises, and the event-driven advantage diminishes (Davies et al., 2018). Mapping conventional dense networks to spiking representations can add conversion overhead and reduce effective sparsity, making comparisons with dense accelerators sensitive to modeling assumptions (Gale et al., 2019; Coleman et al., 2017). As a result, neuromorphic energy wins are most robust in always-on sensing and temporal workloads where sparse events are intrinsic to the data (Merolla et al., 2014).
@@ 2.2. Throughput, latency, and speed
Figure 4. LLM inference bottlenecks and optimization strategies, differentiating between the compute-bound prefill phase and the memory-bandwidth-bound decode phase, and illustrating techniques like KV-cache management, paging, and attention optimization.
For LLM inference, the critical path often alternates between compute-heavy projection/Multilayer Perceptron (MLP) General Matrix Multiplications (GEMMs) and bandwidth-heavy attention/KV-cache reads, as illustrated in Figure 4. Input/Output (IO)-aware attention kernels reduce memory traffic by tiling attention and avoiding materialization of large intermediate tensors, improving both throughput and latency under memory pressure (Dao et al., 2022). Runtime techniques such as paging KV-cache help maintain throughput at high concurrency, but they introduce their own trade-offs in fragmentation, scheduling overhead, and tail latency (Kwon et al., 2023).
At the software level, kernel fusion and code generation reduce launch overhead and improve locality, which matters especially for small GEMMs and latency-sensitive inference. Libraries and compiler stacks (e.g., vendor primitives, graph compilers, and kernel Domain-Specific Languages (DSLs)) play a central role in realizing these optimizations in practice (Chetlur et al., 2014; Chen et al., 2018; Tillet et al., 2019). Consequently, performance claims should specify not only the hardware but also the kernel/library and compilation path used.